The product of two consecutive integers is 19 more than their sum. Find the integers. Write an equation and solve.
Answer by Boreal(15235)

(x)(x+1)=x^2+x
x^2+x-19=2x+1, where x is the smaller integer
x^2-x-20=0
(x-5)(x+4)=0
the integers are 5 and 6, where the product (30) is 19 more than the sum (11)
and
-4, -3, where the product (12) is 19 more than the sum (-7)